A rabbit and a frog jump toward each other.
The rabbit takes two jumps for every three jumps of the frog.
The two jumping animals are 27m away, and it takes the rabbit 4 jumps to cover 1m
How many jumps will the frog take before they are both in same place?

Answer by ikleyn(52814)

.
A rabbit and a frog jump toward each other.
The rabbit takes two jumps for every three jumps of the frog.
The two jumping animals are 27m away, and it takes the rabbit 4 jumps to cover 1m
How many jumps will the frog take before they are both in same place?
~~~~~~~~~~~~~~~~~


Something is missed in this condition, so that the solution is not possible.

Namely, the info regarding the frog rate or his length of jump is missed.


---------------
comment from student:  Thanks.  But the frog rate is obvious in the problem,  right?   frog will jump  6  time for  1  m
---------------


My response.

Would the problem say  "The rabbit makes two jumps at the same time as the frog makes three,  AND
two jumps of the rabbit are of the same length as three jumps of the frog",  then the problem has the simple and easy solution:

    At the same time as the rabbit makes two jumps, the frog makes three, and the distance between them decrease in 0.5 + 0.5 = 1 meter.

    Hence, when the frog will complete his 3*27 = 81-th jumps, the both creatures will be at one place.

Solved.

================

Below is the example of SIMILAR, but ACCURATELY formulated problem from the archive, which I solved at this forum long time ago:

    A rabbit gives 5 jumps at the same time that the dog that chases it gives 4, but 8 dog leaps equate in distance to 11 jumps of the rabbit. 
    If the rabbit takes 66 advantage jumps, how many jumps should the dog give to reach the rabbit? 


The link to this problem in the archive is

https://www.algebra.com/algebra/homework/Percentage-and-ratio-word-problems/Percentage-and-ratio-word-problems.faq.question.1093311.html

https://www.algebra.com/algebra/homework/Percentage-and-ratio-word-problems/Percentage-and-ratio-word-problems.faq.question.1093311.html